OpenStudy (anonymous):

d / (19f^2) / d^4 / (8f^3) How would I go about solving problems like this?

OpenStudy (bibby):

\(\dfrac{\dfrac{d}{19f^2}}{\dfrac{d^4}{8f^3}}\) is that it?

OpenStudy (anonymous):

Yes, that's what it looks like

OpenStudy (bibby):

are you aware that division is multiplication by the reciprocal?

OpenStudy (bibby):

by that, I mean \(\dfrac{3}{4}=3*\dfrac{1}{4}\)

OpenStudy (anonymous):

I've heard that before, I've never really understood it's meaning though haha

OpenStudy (bibby):

basically, you flip the thing you're dividing and multiply. I'll use the draw tool if you want it visually

OpenStudy (anonymous):

The draw tool might help a bit more lol

OpenStudy (bibby):